# De-Hallucinator: Iterative Grounding for LLM-Based Code Completion

### v1

@@ 5.5. RQ4: Efficiency
The following evaluates the efficiency of the approach and how much each of De-Hallucinator’s components contributes to its running time. The preprocessing phase takes, on average, under one second per 1,000 lines of code in a project. For the projects in our dataset, it takes at most 80 seconds, and most projects need at most 26 seconds for the whole preprocessing phase. In a production-level implementation, our CodeQL-based approach could be replaced by using static information that is available on an IDE anyway, which is likely to further reduce the computational effort. Moreover, updating the indexed API references, e.g., when the code base evolves, can be done at low frequency in the background. Retrieving relevant APIs and constructing the augmented prompt takes from 21 to 227 milliseconds per iteration. Finally, on our Nvidia T4 GPUs, the CodeGen v2.5 and the UniXcoder models take, on average, 66.7 seconds and 44.9 seconds for each completion, respectively, and on our Nvidia Tesla V100 GPU, the UniXcoder, CodeGen, CodeGen v2.5, and StarCoder+ models take, on average, 3.6 seconds, 18.6 seconds, 15.3 seconds, and 16.4 seconds per completion, respectively. These numbers are roughly the same for the baseline and for querying the model with De-Hallucinator-augmented prompts. It is important to note that this time is not representative of the time a real-world code completion tool would take. A production-level deployment would run the LLM on a GPU cluster, which typically answers queries within tens to hundreds of milliseconds, as evidenced by the (larger, but not freely available) models of OpenAI.
@@ 6. Limitations and Threats to Validity
One limitation of our work is that we assume an API to be available when completing the code that uses it. However, in some cases, a developer may first write an API usage and then implement the API. In such cases, De-Hallucinator would be unable to retrieve the API reference, and hence, could not improve the completion. To address this limitation, one could configure De-Hallucinator to abstain from repeatedly querying the LLM if the similarity between the initial completion and the retrieved API references is below a threshold. We have implemented and evaluated De-Hallucinator for Python, and although our general approach could be applied to any language, our conclusions are valid only for Python. The set of Python projects we use might not be representative of all Python projects, which we try to mitigate by selecting a diverse set of popular projects. Finally, we have evaluated De-Hallucinator with four LLMs, excluding the models of OpenAI because those models are not publicly available and may even be taken offline at short notice, which would prevent future work from reproducing and comparing against our work. Because these models, like the models we use in our evaluation, are unaware of project-specific APIs, we expect that our approach would also improve their completions.
@@ 7. Related Work
The idea to augment traditional type-based code completion in a data-driven manner was introduced by Bruch et al.(Bruch et al., 2009). More recently, statistical models are used, such as a pre-trained BERT model (Devlin et al., 2018) applied to code completion (Liu et al., 2020), and models trained for specific kinds of completions, e.g., API usages (Raychev et al., 2014) and test methods (Nie et al., 2023). Grammars can help in improving statistical code completions, either by restricting what tokens could be predicted next (Poesia et al., 2022) or by generating code that leaves some syntax subtrees undefined (Guo et al., 2022b). Hellendoorn et al. (Hellendoorn et al., 2019) study data-driven code completion techniques based on recorded real-world completion requests, with a focus on completing a single identifier at a time. Our work differs from all the above by providing project-specific API references as an input to a code completion model.
@@
Motivated by the observation that LLMs lack project-specific information, Shrivastava et al. (Shrivastava et al., 2023) propose a repository-level prompt generation technique to select the best context from a set of predefined contexts to solve the task of line completion. Their method relies on training a separate model that takes a context window around the incomplete line as input, and outputs a ranking for additional contexts. The training routine uses the LLM (in their case Codex) to calculate the loss function. Ding et al.(Ding et al., 2022) describe a similar method, called CoCoMIC, to address the challenge of project-specific APIs. They utilize a custom static analyzer, CCFinder, that initially creates a context graph of program components in the project, and allows retrieval of relevant contexts to complete a statement. They then fine-tuned CodeGen-2B-mono by adding the cross-file contexts to the input. Both of the above are tightly coupled with the underlying LLM: The first approach (Shrivastava et al., 2023) uses the LLM to calculate the loss function for training a new model, and the second approach (Ding et al., 2022) changes the model’s weights during fine-tuning. In contrast, De-Hallucinator queries the LLM as a black-box, and hence, can be easily applied to other models. Moreover, De-Hallucinator introduces a dialog-based way of interacting with an LLM for completion. An approach developed concurrently with ours (Zhang et al., 2023a) includes fragments of project-specific code in the prompt to improve the LLM’s predictions. Similar to our work, they also query the model iteratively. Unlike De-Hallucinator, their approach retrieves existing code fragments, and not API signatures. Since their approach relies on existing code fragments, it can only improve predictions when a project-specific API has already been used before and when this existing usage resembles the desired prediction, whereas our approach applies to all usages of project-specific APIs. Agrawal et al. (Agrawal et al., 2023) modify the decoding stage of the LLM and filter out tokens that are invalid based on static analysis of the existing code. Their approach requires access to the decoder of the LLM, as opposed to our approach which treats the LLM as a black box.
@@
To improve code suggested by LLMs, existing techniques for automated program repair (Le Goues et al., 2019) can be applied in a post-processing step (Fan et al., 2022). Alternatively, the code predicted by a model can serve as input for initializing and guiding a component-based code synthesis algorithm (Rahmani et al., 2021). The above work and ours shares the observation that completions from LLMs often share code elements with the desired code. Instead of improving code in a post-processing step, De-Hallucinator nudges an LLM toward producing better completions by improving the input given to the model.
@@
Work on program repair queries a model multiple times until finding a suitable repair (Lutellier et al., 2020). They repeatedly query the model with the same prompt and may trigger thousands of queries, whereas De-Hallucinator continuously augments the prompt and queries the model only a few times to ensure low latency. Li et al.(Li et al., 2022) propose querying a model with multiple mutations of the given code, and to then use the completion that is closest to the “average” completion. De-Hallucinator differs from their approach by using the initial prediction to construct an improved prompt. Xia et al.(Xia and Zhang, 2023) introduce conversational program repair, which iteratively improves a prompt by adding test failures observed when executing the predicted code. In contrast, we do not require tests or executions, but only information that is statically available in a typical IDE.
@@
The impressive abilities of neural models of code (Pradel and Chandra, 2022) has lead to various other applications beyond code completion. For example, neural models provide type predictions (Hellendoorn et al., 2018; Malik et al., 2019; Pradel et al., 2020; Wei et al., 2020), make predictions about code changes (Hoang et al., 2020; Brody et al., 2020), and enable code search (Gu et al., 2018; Sachdev et al., 2018). LLMs are shown to be useful, e.g., for code mutation, test oracle generation, and test case generation (Bareiß et al., 2022; Schäfer et al., 2023; Kang et al., 2023), and for automated program repair (Jiang et al., 2023).
@@ 8. Conclusion
Motivated by the inability of current LLM-based code completion to correctly predict project-specific APIs, we present De-Hallucinator. Our approach combines the strengths of LLMs, retrieval-based code completion, and iteratively augmenting the prompt with relevant API references. Because De-Hallucinator treats the LLM as a black-box, does not rely on any fine-tuning, and does not require any additional training data, it can be applied to any LLM. Our evaluation on 440 code completion tasks shows that De-Hallucinator significantly improves the quality of code completions over the state-of-the-art baseline, with relative improvements of 23–51% in edit distance and 24–61% in the recall of correctly predicted API usages. Beyond code completion, the idea of using a model’s initial predictions to guide retrieval efforts could be used in other software engineering tasks.
@@
Our implementation and the dataset of incomplete functions is publicly available at https://github.com/AryazE/dehallucinator.

### v3

@@ 5.5. RQ4: Efficiency
The following evaluates the efficiency of the approach and how much each of De-Hallucinator’s components contributes to its running time. The pre-analysis step of code completion using CodeQL takes, on average, under one second per 1,000 lines of code in a project. For the projects in our dataset, it takes at most 80 seconds, and most projects need at most 26 seconds for the whole preprocessing phase. In a production-level implementation, our CodeQL-based approach could be replaced by using static information that is available in an IDE anyway, which is likely to further reduce the computational effort. Moreover, updating the indexed API references, e.g., when the code base evolves, can be done at low frequency in the background. For test generation, the pre-analysis takes between 0.6 seconds and 20 seconds with an average of 3.5 seconds. Because the JavaScript projects in our dataset are smaller than the Python projects, the pre-analysis is faster for test generation.
Retrieving relevant APIs and constructing the augmented prompt for one iteration takes from 21 to 227 milliseconds for code completion, and from 0.1 to 17 milliseconds for test generation. The time to query the LLMs ranges between 1.3 seconds (for the remotely deployed GPT-3.5) and 66.7 seconds (for CodeGen v2.5 running on our local Nvidia T4 GPU), on average per query. These numbers are roughly the same for the baseline and for querying the model with De-Hallucinator-augmented prompts. It is important to note that a production-level deployment would run the LLM on a GPU cluster, which typically answers queries within tens to hundreds of milliseconds, as evidenced by tools like Copilot and Tabnine.
@@ 6. Limitations and Threats to Validity
We assume an API to be available when generating the code that uses it. However, in some cases, a developer may first write an API usage and then implement the API. In such cases, De-Hallucinator would be unable to retrieve the API reference, and hence, could not provide any benefits. To address this limitation, one could configure De-Hallucinator to abstain from repeatedly querying the LLM if the similarity between the initial completion and the retrieved API references is below a threshold. We implement and evaluate De-Hallucinator for Python and JavaScript, and although our general approach could be applied to any language, our conclusions are valid only for these languages. The set of projects we use might not be representative of all projects, which we try to mitigate by selecting a diverse set of popular projects.
@@ 7. Related Work
The idea to augment traditional type-based code completion in a data-driven manner was introduced by Bruch et al.(Bruch et al., 2009). More recently, statistical models are used, such as a pre-trained BERT model (Devlin et al., 2018) applied to code completion (Liu et al., 2020), and models trained for specific kinds of completions, e.g., API usages (Raychev et al., 2014) and test methods (Nie et al., 2023). Grammars can improve statistical code completions, either by restricting the tokens to predict (Poesia et al., 2022) or by generating code that leaves some syntax subtrees undefined (Guo et al., 2022b). Hellendoorn et al. (Hellendoorn et al., 2019) study data-driven code completion based on recorded real-world completion requests, with a focus on completing single identifiers. Our work differs from all the above by providing project-specific API references based on previous completions as an input to a code completion model.
@@
Motivated by the observation that LLMs lack project-specific information, Shrivastava et al. (Shrivastava et al., 2023) propose a repository-level prompt generation technique to select the best context from a set of predefined contexts to solve the task of line completion. Their method relies on training a separate model that takes a context window around the incomplete line as input, and outputs a ranking for additional contexts. The training routine uses the LLM (in their case Codex) to calculate the loss function. Ding et al.(Ding et al., 2022) describe a similar method, called CoCoMIC, to address the challenge of project-specific APIs. They utilize a custom static analyzer, CCFinder, that initially creates a context graph of program components in the project, and allows retrieval of relevant contexts to complete a statement. They then fine-tuned CodeGen-2B-mono by adding the cross-file contexts to the input. Both of the above are tightly coupled with the underlying LLM: The first approach (Shrivastava et al., 2023) uses the LLM to calculate the loss function for training a new model, and the second approach (Ding et al., 2022) changes the model’s weights during fine-tuning. In contrast, De-Hallucinator queries the LLM as a black-box, and hence, can be easily applied to other models.
An approach developed concurrently with ours (Zhang et al., 2023a) includes fragments of project-specific code in the prompt to improve the LLM’s predictions. Similar to our work, they also query the model iteratively. Unlike De-Hallucinator, their approach retrieves existing code fragments, and not API signatures. Since their approach relies on existing code fragments, it can only improve predictions when a project-specific API has already been used before and when this existing usage resembles the desired prediction, whereas our approach applies to all usages of project-specific APIs.
@@
Automated Test Generation
Before the era of LLMs, random feedback-directed test generation became practical through Randoop (Pacheco et al., 2007). LambdaTester (Selakovic et al., 2018) integrates higher order functions into test generation, and Nessie (Arteca et al., 2022) targets asynchronous callbacks. TestPilot (Schäfer et al., 2024) uses LLMs and prompt refinement to generate human-readable regression tests. CodaMosa (Lemieux et al., 2023) uses LLMs to increase the coverage of automatic test generators when stuck in a plateau.
@@
To improve code suggested by LLMs, existing techniques for automated program repair (Le Goues et al., 2019) can be applied in a post-processing step (Fan et al., 2022). Alternatively, the code predicted by a model can serve as input for initializing and guiding component-based code synthesis (Rahmani et al., 2021). The above work and ours shares the observation that completions from LLMs often share code elements with the desired code. Instead of improving code in a post-processing step, De-Hallucinator nudges an LLM toward producing better completions by improving the prompt.
@@
Work on program repair queries a model multiple times until finding a suitable repair (Lutellier et al., 2020). They repeatedly query the model with the same prompt and may trigger thousands of queries, whereas De-Hallucinator continuously augments the prompt and queries the model only a few times. Li et al.(Li et al., 2022) propose querying a model with multiple mutations of the given code, and to then use the completion that is closest to the “average” completion. De-Hallucinator instead uses the initial prediction to construct an improved prompt. Xia et al.(Xia and Zhang, 2023) introduce conversational program repair, which iteratively improves a prompt by adding test failures observed when executing the predicted code. In contrast, we do not require tests or executions, but only information that is statically available in a typical IDE.
@@
The impressive abilities of neural models of code (Pradel and Chandra, 2022) has lead to various other applications beyond code completion. For example, neural models provide type predictions (Hellendoorn et al., 2018; Malik et al., 2019; Pradel et al., 2020; Wei et al., 2020), make predictions about code changes (Hoang et al., 2020; Brody et al., 2020), and enable code search (Gu et al., 2018; Sachdev et al., 2018). LLMs are shown to be useful, e.g., for code mutation, test oracle generation, and test case generation (Bareiß et al., 2022; Schäfer et al., 2024; Kang et al., 2023), and for automated program repair (Jiang et al., 2023).
@@ 8. Conclusion
Motivated by the inability of current LLM-based code generation approaches to correctly predict project-specific APIs, we present De-Hallucinator. The approach exploits the observation that LLMs often predict code that is similar to the desired code, but factually incorrect. We address the hallucination problem by iteratively augmenting the prompt with increasingly relevant API references. Our evaluation on two task, code completion and test generation, shows that De-Hallucinator significantly improves the quality of generations over the state-of-the-art baselines.
@@
Our implementation, datasets, and evaluation scripts are publicly available at https://github.com/AryazE/dehallucinator and https://github.com/AryazE/testpilot.